In studying his campaign plans, Mr. Singleton wishes to estimate the difference between men's and women's views regarding his appeal as a candidate. He asks his campaign manager to take two random independent samples and find the 95% 95 % confidence interval for the difference. A random sample of 559 559 male voters and 504 504 female voters was taken. 173 173 men and 210 210 women favored Mr. Singleton as a candidate. Find this confidence interval. Step 2 of 4 : Find the critical value that should be used in constructing the confidence interval.
Answer #1

Solution:

Solution :

Given that,

n1 = 559

x1 = 173

n2 =504

x2 = 210

Construct 95% confidence interval for the difference between proportions.

Step 2 of 4 : Find the critical value that should be used in constructing the confidence interval.

Our aim is to construct 95% confidence interval.

c = 0.95

= 1- c = 1- 0.95 = 0.05

  /2 = 0.05 2 = 0.025 and 1- /2 = 0.975

Search the probability 0.975 in the Z table and see corresponding z value

= 1.96   

Critical value is 1.96